# SLIM-Brain: A Data- and Training-Efficient Foundation Model for fMRI Data Analysis

### v2

@@ Abstract
Foundation models are emerging as a powerful paradigm for fMRI analysis, but current approaches face a dual bottleneck of data- and training-efficiency. Atlas-based methods aggregate voxel signals into fixed regions of interest, reducing data dimensionality but discarding fine-grained spatial details, and requiring extremely large cohorts to train effectively as general-purpose foundation models. Atlas-free methods, on the other hand, operate directly on voxel-level information - preserving spatial fidelity but are prohibitively memory- and compute-intensive, making large-scale pre-training infeasible. We introduce SLIM-Brain (Sample-efficient, Low-memory fMRI Foundation Model for Human Brain), a new atlas-free foundation model that simultaneously improves both data- and training-efficiency. SLIM-Brain adopts a two-stage adaptive design: (i) a lightweight temporal extractor captures global context across full sequences and ranks data windows by saliency, and (ii) a 4D hierarchical encoder (Hiera-JEPA) learns fine-grained voxel-level representations only from the top-$k$ selected windows, while deleting about 70% masked patches. Extensive experiments across seven public benchmarks show that SLIM-Brain establishes new state-of-the-art performance on diverse tasks, while requiring only 4 thousand pre-training sessions and approximately 30% of GPU memory comparing to traditional voxel-level methods.
@@ 4.1 Dataset
We leverage self-supervised pretraining on four public neuroimaging datasets: a single HCP session (Van Essen et al., 2013), CHCP (Ge et al., 2023), and two releases from the Amsterdam Open MRI Collection (AOMIC)—PIOP1 and PIOP2 (Snoek et al., 2021); the Adolescent Brain Cognitive Development (ABCD) Study Casey et al. (2018). We use 70% of the data (4129 sessions) for training, with the remainder reserved for validation (10%) and testing (20%). External validation spans four datasets: the Autism Brain Imaging Data Exchange (ABIDE) (Di Martino et al., 2014); the ADHD-200 Sample (ADHD) (consortium, 2012); Alzheimer’s Disease Neuroimaging Initiative (ADNI) Jack Jr et al. (2008); Parkinson’s Progression Markers Initiative (PPMI) Marek et al. (2011). See Appendix  A.1 and  A.2 for preprocessing and dataset details.
@@ 4.2 Fine-tuning results
We fine-tune our model and evaluate it on out-of-distribution datasets, as summarized in Table 1 and Appendix A.3. We benchmark against seven representative fMRI models—BrainNetCNN (Kawahara et al., 2017), BrainGNN Li et al. (2021), BrainLM (Caro et al., 2023), BrainMass (Yang et al., 2024), Brain-JEPA (Dong et al., 2024), NeuroSTORM Li et al. (2025) and SwiFT Kim et al. (2023)—using the authors’ released codes or checkpoints (pretrained on 1010k–6464k samples). Note that BrainMass includes these datasets in its pre-training corpus, whereas our method utilizes them strictly for fine-tuning. We adopted the medium-sized model (45M parameters) trained on 4,129 sessions as the default configuration for the experiments. Across all the OOD evaluations, our model consistently outperforms the baselines, indicating that it learns domain-invariant representations that transfer to unseen datasets and tasks with minimal adaptation. Task description (A.2), statistical method (A.3), internal experiment results (A.4), linear probing results results (A.5), configuration details for ours (A.6) and baselines (A.7) are provided in the Appendix.
@@ 4.3 Ablation Study
Table 2: Table B: Ablation study on key frame selection strategies (ADHD). We report Mean and STD over 3 independent runs. The symbol * indicates statistical significance ($p<0.05$).
@@ A.1 Preprocessing
A.2 Dataset and Downstream Tasks
Human Connectome Project (HCP).
We utilize the HCP dataset as our primary benchmark for pre-training and initial evaluation, comprising 1,011 subjects. This cohort represents a young adult population with ages ranging from 22 to over 36 years. Specifically, the age distribution includes distinct groups: 22–25 ($N=218$), 26–30 ($N=437$), 31–35 ($N=346$), and 36+ ($N=10$). The dataset maintains a balanced sex ratio with 469 males and 542 females, providing high-quality, standardized data ideal for establishing baseline model performance. We only use one session per participants Van Essen et al. (2013). HCP Fingerprint Fingerprint serves as a subject-identification benchmark. We randomly chose 100 subjects as distinct classes. Three different sessions from the same subjects are assigned to train, validation, and test splits, ensuring that the task evaluates cross-session fingerprint consistency rather than session-specific artifacts.
Amsterdam Open MRI Collection (AOMIC).
We include two independent datasets from the AOMIC project for pretraining. PIOP1 consists of 203 subjects with a mean age of $22.208\pm 1.795$ years (86 males, 117 females). PIOP2 comprises 222 subjects with a similar mean age of $21.946\pm 1.783$ years (95 males, 127 females).
Chinese Human Connectome Project (CHCP).
To enhance the demographic diversity of our pre-training corpus, we incorporate the CHCP dataset, which consists of 304 subjects (145 males, 159 females).The CHCP cohort presents a significantly broader age distribution ($29.326\pm 14.860$ years).
Adolescent Brain Cognitive Development (ABCD).
To evaluate model generalization on a larger scale and a distinct demographic, we incorporate the ABCD dataset, which includes 2,398 subjects Casey et al. (2018). In contrast to HCP, this cohort focuses on early brain development, with a tightly clustered mean age of $9.925\pm 0.627$ years. And the dataset features a highly balanced sex distribution (1,226 males and 1,172 females).
ADHD Disease
ADHD-200 is an open multi-site neuroimaging resource with resting-state fMRI (rs-fMRI) across multiple contributing centers consortium (2012). For this task we merge all ADHD subtypes into a single ADHD label and it contains 938 participants (ADHD 356 vs. control 582).
Autism Brain Imaging Data Exchange (ABIDE)
The Autism Brain Imaging Data Exchange is a multi-site, open-access initiative that aggregates structural and resting-state fMRI scans - alongside rich phenotypic data - from individuals with autism spectrum disorder (ASD) and matched typically developing controls to accelerate reproducible neuroimaging research Di Martino et al. (2014). It contains 871 participants (mean age 16.9, std 7.5). We perform regression on age and scale both the labels and results to the range [0, 1] using StandardScaler. Additionally, we divide the age variable into four balanced groups for the classification task.
Alzheimer’s Disease Neuroimaging Initiative (ADNI)
The Alzheimer’s Disease Neuroimaging Initiative is a large‐scale, longitudinal, multi‐center project initiated in 2004, designed to collect and publicly share clinical data from cognitively normal older adults, individuals with mild cognitive impairment (MCI) and patients with Alzheimer’s disease (AD) Jack Jr et al. (2008). For our task, we separated disorders into two classification with healthy controls.
Parkinson’s Progression Markers Initiative (PPMI)
The Parkinson’s Progression Markers Initiative is an ongoing, international, longitudinal observational study launched in 2010 by The Michael J. Fox Foundation to identify and validate biomarkers of Parkinson’s disease (PD) risk, onset, and progression. We combined two different diseases and healthy controls from the PPMI dataset to create a three-class classification task Marek et al. (2011). This cohort consists of 474 subjects (272 males, 202 females; mean age 66.44). We utilize the data for a three-class disease classification task: Prodromal ($303$), PD ($142$), and Control ($29$).
A.3 Statistical analysis of external tasks.
@@ A.4 Internal experiments
A.5 Linear probing Results
@@ A.6 Experimental settings
A.7 Details of baseline models
In this section, we introduce our baselines models. First four foundational models for brain activity prediction: BrainLM, BrainMASS, Brain-JEPA and NeuroSTORM, each of which uses distinct parcellation schemes and is trained on large-scale, multi-site fMRI datasets. We also fine-tuned our downstream tasks using other end-to-end models: BrainNetCNN, BrainGNN, SwiFT.
BrainNetCNN is a convolutional neural network specifically designed for analyzing connectome-structured data. Unlike traditional models that process images or graphs, BrainNetCNN employs specialized convolutional filters—edge-to-edge, edge-to-node, and node-to-graph—that directly model pairwise connectivity patterns. In our experiments, the model takes as input a functional connectivity (FC) matrix, which is computed using Pearson correlation coefficients between regional fMRI time series. We set the hyperparameters as follows: batch size of 64, learning rate of 0.01, and weight decay of 0.05.
BrainGNN is a graph neural network architecture designed to analyze brain connectomes at the population level by learning interpretable node- and region-specific biomarkers. For each subject, it construct a graph whose nodes correspond to brain regions and whose edges encode statistical dependencies between regional fMRI time series. The functional connectivity (FC) matrix is used as the node-level feature representation, and subject-level labels provide supervision for learning anatomically meaningful subgraphs and node embeddings. Partial correlations between fMRI time series are computed and assigned as edge attributes, enabling the model to capture conditional dependencies between brain regions rather than only marginal pairwise correlations. During training, we use batch size of 64, learning rate of 0.001, and weight decay of 0.5. It is important to note that, unlike the multiple regularization terms used in classification tasks, such as entropy regularization and consistency regularization, we only utilize the Mean Squared Error (MSE) loss for the age regression task to minimize adaptations. However, this approach has led to unstable learning in the BrainGNN model, resulting in its failure to converge on this task, even after extensive hyperparameter tuning.
BrainLM is the first fMRI foundation model specifically designed to capture the spatiotemporal dynamics of brain activity through a Transformer-based masked autoencoder architecture. It employs the AAL-424 atlas for the parcellation of brain regions, transforming fMRI recordings into a 424-dimensional set of regions of interest (ROIs). The model is trained on an extensive dataset consisting of 6,700 hours of fMRI data derived from 77,298 recordings across the UK Biobank and the Human Connectome Project. For downstream fine-tuning, we choose the following hyperparameters: batch size of 64, learning rate of 0.00001, and training duration of 30 epochs.
BrainMASS utilizes the Schaefer 100-region atlas for the parcellation of brain networks. It was trained on 26 datasets, encompassing a total of 64,584 subjects, including data from the UK Biobank (UKB), Human Connectome Project (HCP), and ADHD-200, etc. The model is designed to learn representations of functional brain networks by integrating both masked ROI modeling and latent representation alignment techniques, thereby enhancing its diagnostic performance. The following hyperparameters were selected for downstream fine-tune: a batch size of 64, a learning rate of 0.0002, and a training duration of 100 epochs.
Brain-JEPA employs an advanced parcellation strategy that combines the Schaefer-400 cortical regions with the Tian-Scale III subcortical regions, resulting in a total of 450 regions of interest (ROIs). The model is trained on 80 % of the UK Biobank (UKB) data and utilizes a joint-embedding predictive architecture. This architecture incorporates both spatial and temporal masking techniques to capture dynamic patterns of brain activity. The model is trained with the following hyperparameters: a batch size of 16, a learning rate of $4\times 10^{-4}$, over a total of 50 epochs.
NeuroSTORM establishes a general-purpose neuroimaging foundation model that learns directly from raw 4D fMRI volumes, avoiding the information loss associated with projecting data onto pre-defined atlases or connectomes. It employs a Shifted-Window Mamba (SWM) backbone. The model is pre-trained on a large-scale fMRI data from over 50,000 subjects—spanning the UK Biobank, ABCD, and HCP datasets. The model utilizes a Spatiotemporal Redundancy Dropout (STRD) module to mitigate data redundancy while capturing long-range dependencies.
SwiFT (Swin 4D fMRI Transformer) addresses the challenge of modeling high-dimensional spatiotemporal brain dynamics by learning directly from raw 4D fMRI volumes, thereby avoiding the information loss introduced by hand-crafted features. Through its 4D windowed attention mechanism and computationally efficient architecture, SwiFT outperforms recent state-of-the-art models on large-scale datasets. Moreover, contrastive self-supervised pretraining further enhances its downstream performance, highlighting SwiFT’s effectiveness as an end-to-end framework for functional brain imaging. In our experiments, we fine-tune SwiFT with a batch size of 16, a learning rate of $1\times 10^{-6}$, a weight decay of 0.01, and a training duration of 9 epochs. Due to SwiFT’s substantial I/O demands, we run each downstream task only once.
A.8 Details of key frame selection
@@
A.9 Details of scaling study
@@
A.10 Detailed implementations of our ablation structures
@@
A.11 Neurosynth& Integrated gradients

### v3

@@ Abstract
Foundation models are emerging as a powerful paradigm for fMRI analysis, but current approaches face a dual bottleneck of data- and training-efficiency. Atlas-based methods aggregate voxel signals into fixed regions of interest, reducing data dimensionality but discarding fine-grained spatial details, and requiring extremely large cohorts to train effectively as general-purpose foundation models. Atlas-free methods, on the other hand, operate directly on voxel-level information - preserving spatial fidelity but are prohibitively memory- and compute-intensive, making large-scale pre-training infeasible. We introduce SLIM-Brain (Sample-efficient, Low-memory fMRI Foundation Model for Human Brain), a new atlas-free foundation model that simultaneously improves both data- and training-efficiency. SLIM-Brain adopts a two-stage adaptive design: (i) a lightweight temporal extractor captures global context across full sequences and ranks data windows by saliency, and (ii) a 4D hierarchical encoder (Hiera-JEPA) learns fine-grained voxel-level representations only from the top-$k$ selected windows, while deleting about 70% masked patches. Extensive experiments across seven public benchmarks show that SLIM-Brain establishes new state-of-the-art performance on diverse tasks, while requiring only 4 thousand pre-training sessions and approximately 30% of GPU memory comparing to traditional voxel-level methods.Code and trained weights of SLIM-Brain are available at link.
@@ 4.1 Dataset
We leverage self-supervised pretraining on four public neuroimaging datasets: a single HCP session (Van Essen et al., 2013), CHCP (Ge et al., 2023), and two releases from the Amsterdam Open MRI Collection (AOMIC)—PIOP1 and PIOP2 (Snoek et al., 2021); the Adolescent Brain Cognitive Development (ABCD) Study Casey et al. (2018). We use 70% of the data (4129 sessions) for training, with the remainder reserved for validation (10%) and testing (20%). External validation spans four datasets: the Autism Brain Imaging Data Exchange (ABIDE) (Di Martino et al., 2014); the ADHD-200 Sample (ADHD) (consortium, 2012); Alzheimer’s Disease Neuroimaging Initiative (ADNI) Jack Jr et al. (2008); Parkinson’s Progression Markers Initiative (PPMI) Marek et al. (2011). See Appendix  A.1 for preprocessing details.
@@ 4.2 Fine-tuning results
We fine-tune our model and evaluate it on out-of-distribution datasets, as summarized in Table 1 and Appendix A.2. We benchmark against seven representative fMRI models—BrainNetCNN (Kawahara et al., 2017), BrainGNN Li et al. (2021), BrainLM (Caro et al., 2023), BrainMass (Yang et al., 2024), Brain-JEPA (Dong et al., 2024), NeuroSTORM Li et al. (2025) and SwiFT Kim et al. (2023)—using the authors’ released codes or checkpoints (pretrained on 1010k–6464k samples). Note that BrainMass includes these datasets in its pre-training corpus, whereas our method utilizes them strictly for fine-tuning. We adopted the medium-sized model (45M parameters) trained on 4,129 sessions as the default configuration for the experiments. Across all the OOD evaluations, our model consistently outperforms the baselines, indicating that it learns domain-invariant representations that transfer to unseen datasets and tasks with minimal adaptation. Statistical method (A.2), internal experiment results (A.3), linear probing results results (A.4), configuration details for ours (A.5) are provided in the Appendix.
@@ 4.3 Ablation Study
Table 2: Ablation study on key frame selection strategies (ADHD). We report Mean and STD over 3 independent runs. The symbol * indicates statistical significance ($p<0.05$).
@@ A.1 Preprocessing
A.2 Statistical analysis of external tasks.
@@ A.3 Internal experiments
A.4 Linear probing Results
@@ A.5 Experimental settings
A.6 Details of key frame selection
@@
A.7 Details of scaling study
@@
A.8 Detailed implementations of our ablation structures
@@
A.9 Neurosynth& Integrated gradients